Four equal squares are placed side by side to make a rectangle whose perimeter is 15 meters. Find the area of the rectangle in square meters? ( answer is 9 - show your solution )
Answer by ptaylor(2198)

Perimeter of a rectangle(P)equals two times Length(L) plus two times Width(W) or P=2L+2W
Let S=length of the side of the squares
Then the length of the rectangle is 4S
And the width of the rectangle is S
Sooo
Perimeter of rectangle=15 meters=2(4S)+2S or
8S+2S=15
10S=15
S=15/10=(3/2) meters
Area of rectangle(A)=Length(L) times Width(W) or A=LW
Length of rectangle=4(3/2)=6 meters
Width of rectangle=3/2 meters
Area of of rectangle(A)=6*(3/2)=9 sq meters

Hope this helps---ptaylor